A firm in a perfectly competitive market has the following cost function: C = 1/3q^3 - 5q^2 + 3q + 10. If the market clearing price is 6, obtain the profit maximising level of output.
1
Expert's answer
2020-10-13T10:45:05-0400

solution

cost function

c(q)="\frac{1}{3}q^3-5q^2+3q+10"

MCP=6


for maximum profit will be at the point where


P=MCP=MC


and

MC="\frac{dC(q)}{dq}"

"MC=\frac{d}{dq}(\frac{1}{3}q^3-5q^2+3q+10)"


MC="q^2-10q+3"


and MPC=6

so

"q^2-10q+3=6"


"q^2-10q-3=0"

by solving it

"q=-0.29\space\space ,\space10.29"


by ignoring negative value , the profit maximising level of output will be q=10.29.